OpenStudy (anonymous):

solve 18 divided by 3 times 3 - 2 times 3/ 2 times 3 - 2

OpenStudy (anonymous):

18/3*3-2*3*3-2 6*3-2*3*3-2 18-2*3*3-2 18-6*3-2 18-18-2 -2

OpenStudy (anonymous):

you wrote the problem wrong... Look at the way i wrote it=)

OpenStudy (anonymous):

18/3*3-2*3/2*3-2 6*3-2*3/2*3-2 18-2*3/2*3-2 18-6/2*3-2 3/2*3-2 3/6-2 3/4 =3/4 or 0.75 sorry about that.

OpenStudy (anonymous):

uhhhh thats not one of the multiple choice options...... my teacher actually told me the answer was 3 but im trying to figure out how to get it....any ideas????

OpenStudy (anonymous):

Starting from 18-6/2*3-2 I messed up again sorry 12/2*3-2 12/6-2 12/4 =3 THERE! :D

OpenStudy (anonymous):

wait can u explain the first step you did

OpenStudy (anonymous):

18/3*3-2*3/2*3-2 6*3-2*3/2*3-2 18-2*3/2*3-2 18-6/2*3-2 12/2*3-2 12/6-2 12/4 =3

OpenStudy (anonymous):

when you do 6 times 3 - 2 times 3... where do you get the 6 from?

OpenStudy (anonymous):

from 18/3.